Question 21 The ka for hypobromous acid, HBrO, is 2.0 x 109. What is the pH of a 0.800 M KBro solution at 25°C? NA tion 3
Answer #1


Concentration of KBrO ion is 0.800M Ka of HB10 is 2.0 x10-9 Kb= 1.0*10-24 2.0x100 =5.0*10-6 Hydrolyis of the potassium format